OpenStudy (anonymous):

Let f(x) = 7x - 13. Find f-1(x). this stuff is really confusing me plz help

OpenStudy (anonymous):

f(x) = y = 7x - 13 you should change it to x=something!

OpenStudy (anonymous):

how would i do that im just starting this stuff for my sofmore year so im really new at it

OpenStudy (anonymous):

y = 7x - 13 y + 13 = 7x (y + 13)/7 = x f^-1(x) = (x + 13) / 7 it's clear or need explain? ;)

OpenStudy (anonymous):

so i would solve (x+13)/7?

OpenStudy (anonymous):

do you agree with this \[f(a) = b \rightarrow f^-1(b) = a\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

now\[f(x) = 7x - 13\]and we know\[f ^{-1}(7x - 13) = x\]now for found f inverse we do this\[f(x) = y = 7x - 13 \rightarrow y+13 = 7x \rightarrow \frac{ y+13 }{ 7 }=x\]now just relocation x and y\[\frac{ x+13 }{ 7 }=y=f ^{-1}(x)\]